ULTRAFAST PHOTOPHYSICS OF SIMPLE AROMATIC CHROMOPHORES

RAUL MONTERO, ALVARO PERALTA CONDE, MARTA FERNÁNDEZ-FERNÁNDEZ, FER-NANDO CASTANO and \underline{ASIER LONGARTE}, Departamento de Química Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco (UPV-EHU), Ap. 644, E-48080 Bilbao, Spain.

It is known that the photophysical properties of chromophores is decisively conditioned by the existence of πσ^{∗} transitions with repulsive character, in the vicinity of the characteristic bright ππ^{∗} absorptions. The interplay between these two types of states triggers a complex dynamics that takes place in the femto-picoseconds scale.^{a} Our group is interested in understanding the ultrafast relaxation mechanisms involving πσ^{∗} surfaces in a set of isolated aromatic chromofores including: aniline, indole, phenol and pyrrole. In the research carried out, the dynamical signature of the πσ^{∗} states has been tracked directly on the parent molecules photoexcited in a broad range of their near UV absorption specta, by multiphoton delayed ionization at several wavelengths. For the studied molecules, the work provides a detailed view on the relaxation pathways available, while permits to relate their photophysical behavior with the relative location of the πσ^{∗} and ππ^{∗} states.

aW. . Domcke et. al., Science, \underline{302}, 1693, 2003.